OpenStudy (anonymous):

How do I simplify 282/329

OpenStudy (anonymous):

329 is 7*47...

OpenStudy (aroub):

=6/7

OpenStudy (anonymous):

FIRST FIND THE FACTORS OF 282 SINCE 282 IS AN EVEN NUMBER, IT SEEMS EASIER. 282 = 2 * 141. Now, quickly add up the digits in 141, you get 6, then you know 141 has at least a factor 3. So you further divide 141 by 3 to get 47. Now, 47 is a prime number, so you may write here 282 = 2 *3*47. Next, check to see if any of the factors of 282 can be divided into 329. 329 is an odd number, but is not divisible by 3 because all three digits add up to be 14, not divisible by 3, so the only candidate is 47. Take 329 and divide it by 47 to get a nice whole number 7. Therefore, to simplify 282/329, the common factor for 282 and 329 is 47, you divide both the numerator and the denominator by 47 to get 6/7. This is the answer.